40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 9 00:32:39 2013, rana, Configuration, IOO, Turn on MCL 
    Reply  Wed Apr 10 03:59:41 2013, Den, Configuration, IOO, Turn on MCL MCL.pdf
Message ID: 8434     Entry time: Wed Apr 10 03:59:41 2013     In reply to: 8426
Author: Den 
Type: Configuration 
Category: IOO 
Subject: Turn on MCL 

Quote:

 My belief is that the frequency noise from the unstabilized MC is making the PRC locking harder. This will be investigated by tuning the shape of the MCL/MCF crossover so that we can turn it on without ruining the arm cavity spectra. Since the PRC length is ~2x smaller than the MC, we would expect it to be less sensitive to the MC frequency noise. But, since there is some common mode rejection in there, this may not be true. We'll only know by measuring PRC control signal with MCL on/off.

 I think if we make MCL UGF higher then 20 Hz, arm cavity spectra will feel it. It might be possible to use a combination of feedback and feedforward control from ground seismometers. I made MCL UGF at 3 Hz to reduce 1 Hz motion of the pendulum; feedforward OAF subtracted the stack at 3.3 Hz. Once OAF converged, I blocked adaptation and the filter became static FIR. MC length RMS was reduced by a factor of 10 and arm cavity spectra was not affected at frequencies >20 and became better at low frequencies. We'll see if this enough.

On the attached plot red color shows MC_F with MC_L OFF, blue - MC_L is ON, green - MC_L and OAF are ON.

Then I locked PRCL (using AS_Q and REFL55_I) to carrier and aligned the cavity. Power RIN was 50-70% and 00 beam on the POP camera was moving significantly. BS oplev was shaking the optics at 5 Hz. I fixed it, but there should be something else as RIN was still high.

Attachment 1: MCL.pdf  30 kB  | Hide | Hide all
MCL.pdf
ELOG V3.1.3-